if the sum of the squares of the distances of a point from the three coordinate axis be 36, then its distance from the origin is

Grade:11

1 Answers

Mustafa Ahmed
7 years ago
Let the distances be x y and z . Given that X^2+Y^2+Z^2=36. By distance formula we have: distance of a point from origin is the square root of the sum of squares ofndistances of that point along the axes. Mathematically, its root (X^2+Y^2+Z^2). = root (36) = 6 or -6. Since distance cant be negative, -6 neglected hence 6 is answer.